# A Set-Sequence Model for Time Series

### v1

@@ 5.2 Generalization Across Number of Input Units
Task
Given the setting with $M$ units as described above, the task is to predict the next state transitions accurately based on the previously seen data. We observe $\bar{N}_{t}$ which is the fraction of units defaulted when only $n$ out of $M$ units are observed, hence a noisy realization of $N_{t}$. In the default prediction setting, this correspond to having a large universe of loans evolving jointly, but only observing a limited subset of the data.
Baseline Based on the Kalman Filter
We develop a strong baseline that represents near ideal performance on the task when trained with infinite data. We assume that the true transition matrix is known, and we also assume that the dynamics of $\lambda_{t}$, $\lambda_{t+1}=\beta\lambda_{t}+\alpha N_{t}$ are known, with the only unknown factor being not observing the true fraction of transition to the absorbing state $N_{t}$, but we only observe $\bar{N}_{t}$, where $\bar{N}_{t}$ is the observed fraction given n out of M units. The problem is now to infer the best estimate of $\lambda_{t}$ based on partial observations $\bar{N}_{1},\dots,\bar{N}_{t}$. In Appendix D.3 we derive an estimation $\hat{\lambda}_{t}$ that is near optimal, based on a Kalman Filter. This estimate is then used along with the true transition dynamics to generate transition probabilities for the next time period for each unit. The goal is hence to achieve performance as close as possible to the Kalman Filter baseline.
Model Training
During training we draw the number of units per batch as
$D\sim(1-\gamma)\cdot\delta_{M}+\gamma\cdot\text{LogUniform}(1,M),$
with $\gamma=0.08$; Thus most batches use all $M=1000$ units while a few expose smaller pools for generalization. An analysis of other sampling schemes along with more training details are provided in Appendix D.
Metrics
We evaluate models on their ability to predict transitions to the absorbing (rare) state using three metrics: AUC, correlation, and $R^{2}$. The AUC measures the model’s ability to rank true absorbing transitions above non-absorbing ones. Correlation captures the linear agreement between predicted and true transition probabilities. $R^{2}$ quantifies the variance explained for the absorbing transitions specifically. See Appendix 5.2 for formal definitions.
Results
@@
Figure 2: Comparison between the Set-Sequence model, MHA-Sequence model, and the Kalman Filter, by considering the AUC, $R^{2}$, and Correlation, each for the absorbing (rare) state. The Set-Sequence model reaches near the oracle-like Kalman Filter model across the full range of units for inference, for all the considered metrics.
In Figure 2, the performance of both the Set-Sequence and the MHA-Seq model are compared with the oracle-like Kalman Filter baseline for inference sizes ranging from one unit (out of 1000) to the full information setting with 1000 units. We note that both the Set-Sequence and the MHA-Seq models reach near optimal performance on the task for all considered metric, across the full spectrum of number of input units. This shows that a single model generalizes at inference to unseen or rarely seen input sizes from 1 to 1000 units. An important consideration is that the oracle-like Kalman Filter baseline are computed using the infinite data limit, and the set sequence model is trained on a restricted size training set. The MHA-Seq offers a slight performance advantage on the synthetic task, but this comes at a computational disadvantage due to quadratic scaling in the number of input units.
Interpretability
The Set-Sequence model produces interpretable results. Table 3 shows the average correlation across 100 samples of length 100 from the test set, between the first set variable in layer i of the Set-Sequence and the latent factor $\lambda_{0,t}$. As we observe more units the correlation increases, with a very high average correlation $0.951$ in the fully observed case in layer 3. Appendix D.2 contains visualizations of the learned set variables.
@@
Table 3: Average correlation for 100 samples, each with 100 time steps between the set summary at layer i and the true latent variable, $\lambda_{0,t}$. Best values in each column are in bold, with the second best underlined.
@@
6 Empirical Application: Equity Portfolio Prediction
We evaluate the Set-Sequence model on a portfolio prediction task. In the task, given a universe of $N$ assets, at the current time, the model outputs $N$ portfolio weights, indicating the portfolio we invest in. The portfolio returns are realized from the inner product between the next day returns and the chosen portfolio weights. The goal is to optimize the Sharpe ratio, a measure of the risk-adjusted returns. This is a common objective in the asset pricing literature, as in [11].
@@ 6 Empirical Application: Equity Portfolio Prediction
We use data from the Center for Research in Security Prices (CRSP). The full dataset consists of 36600 stocks. For each period, we train on 8 years and test the model on the following year. For each model fitting, we filter out stocks where any returns in the train or test period are not available. After this filter, we filter to include only the 500 assets with the largest market cap at the last quarter of the train period, mimicking the S&P 500 index. This ensures that we consider assets that are liquid with small bid-ask spreads, aligned with prior work [11]. We use a subset of the features from [5], with the same normalization (cross-sectional rank normalization). The features are based on recent cumulated returns, volatility, volume, and quarterly firm characteristics (such as book-market ratio, and total assets). In total we use 79 features listed in Table 10 and Table 11 in the Appendix.
@@
Let $X^{i}_{t}$ be the covariates at time t, and let $Y^{i}_{t}$ be the excess daily returns for asset i (returns with the risk free rate subtracted), where $X_{t}$ is the covariates for all assets, and $Y_{t}$ is the returns for all assets. We then use $X_{1},\dots,X_{t-1}$ to predict $\hat{w}_{t}$, the portfolio weights at the next step. Given a series of daily excess returns, $r_{1},\dots,r_{T}$, where $r_{t}=Y_{t}^{T}\hat{w}_{t}$, the daily Sharpe ratio (SR) is then
$SR=\frac{\bar{r}}{\sqrt{\frac{1}{T}\sum_{t=1}^{T}(r_{t}-\bar{r})^{2}}},$ (6)
where $\bar{r}=\frac{1}{T}\sum_{t=1}^{T}r_{t}$. We $L_{1}$ normalize the portfolio weights at each timestep to maintain unit leverage ($||w_{t}||_{1}=1$). Aligned with [11] we use the negative of the Sharpe ratio as the optimization objective.
Baselines
We benchmark against the sequence baselines described in Section 4. Each baseline is trained on a single unit at a time with weights shared across units—a strategy that generalizes far better than feeding the entire cross-section, which performs poorly (see Section 5.1). We further compare with task-specific models: the CNN-Transformer and the additional baselines of [11], which use aligned dataset, time span, and feature set (the latter following [5]).
Metrics
The main metric to evaluate the performance is the annualized Sharpe ratio. We also give the percentage yearly return, the yearly return volatility, the average daily turnover, defined as the average $L_{1}$ distance between the weights in consecutive days. Beta is the covariance of portfolio and market returns divided by the market variance, measuring market exposure. A low beta is desirable to ensure returns are independent of market direction.
Results
Table 4 shows the aggregate performance. The Set-Sequence model outperforms every sequence model baseline. We note a high annualized mean Sharpe ratio over the full period of 4.82, 22% higher than the second best sequence model (S4), and 32% higher than the LongConv model, which the Set-Sequence model use for the sequence component in the experiment. Compared with the other sequence models, the Set-Sequence model shows robust results, with a Sharpe ratio standard deviation over 5 random initialized training runs of 0.12, compared with 0.52 for the Transformer and 0.29 for S4.
Figure 3 shows the cumulative returns for the models. We also compare the Set-Sequence model with domain specific models on the time period Jan. 2002 - Dec. 2016 to be aligned with time period of prior work. Table 5 reports that the general Set-Sequence model outperforms the CNN-Transformer designed specifically for the equities task by 42 % in terms of the Sharpe ratio. In Section E in the Appendix we include an analysis of the Sharpe ratio when we include transaction costs. We use a net Sharpe objective function and show that the Set-Sequence model still outperforms all baselines in the presence of transaction costs.
[…]
@@ Results
[…]
Set-Seq (Ours) $\textbf{4.82}\,\pm\,0.12$ 13.0 2.69 0.028 0.91 0.48
Table 5: Out-of-sample annualized performance metrics from 2002 to 2016. All CNN-Transformer models are based on the IPCA method described in [11]. For the Set-Sequence model the results are the mean over 5 training initialization seeds. For the Sharpe we also report its standard deviation over the seeds.
@@
7 Empirical Application: Mortgage Risk Prediction
Next we evaluate the Set-Sequence model on a mortgage risk prediction task, with the goal to predict the state (current, 30 days delinquent, 60 day delinquent, 90+ days delinquent, Foreclosure, Paid Off, and Real Estate Owned) in the next month given a history of 52 covariates (FICO score, loan balance, current interest rate etc.) for a mortgage.
Dataset
We source the data from 4 ZIP codes in the greater LA area as a subset of the CoreLogic Loan-Level Market Analytics Dataset, as these are among the ZIP codes with most active mortgages. We follow the same data filtering procedure as in [22], resulting in 5 million loan month transitions from 117523 active loans. We use the following train/validation/test split: Jan. 1994 - Jun. 2009/Jul. 2009 - Dec. 2009/Jan. 2010 - Dec. 2022. The number of active loans in the dataset varies between 4000 in Jan 1994, up to 22000 in 2006. In Appendix F.1, we include additional dataset descriptions, filtering procedure, transition probabilities and counts of the number of active loans over time.
Figure 4: AUC gain (Set-Seq – baseline [22]) for each transition. Deeper green indicates a larger Set-Seq advantage; red indicates the baseline performed better. Larger circles indicate more common transitions, with size proportional to the log of their frequency. Only transitions occurring at least 10 times are included.
Setup
We compare the Set-Sequence model with the domain specific models presented in [22], representing the current state of the art, along with the sequence models from Section 4. The models from [22] includes a neural network baseline has 5 layers with early stopping and dropout regularization, and a logistic regression baseline. Both baselines use the features at the current time to predict the state in the following month. For each training batch, for a given time we sample 2500 loans out of all active ones. We use a sequence length of 50, and use a multi-class cross entropy loss during training.
Results
Table 6: Comparing performance of different models on the mortgage risk prediction task on the test set. The average AUC is the mean of the AUCs to go between different states. The average transition probabilities are the average probability for the correct class over the full test set.
@@
Set-Sequence (Ours) $\mathbf{0.200}$ $\mathbf{0.683}$
In Figure 5 we see that the Set-Sequence model robustly outperforms the prior state of the art model on the task, with better AUC for 22 out of 25 transitions, including the most common ones and economically important transitions, such as those to the paid-off state from current. In Figure 5, the foreclosure rate, a known source of cross unit dependency, is shown to have a Pearson correlation of 0.67 with the set summary in the first Set-Sequence layer, indicating the interpretability of the learned set summaries. Table 6 shows that the Set-Sequence model achieves the best performance on both cross-entropy loss and average AUC, outperforming all domain-specific and general sequence model baselines. In particular, it improves average AUC by 4 points over the best domain-specific baseline (5-layer neural network).
@@ 8 Conclusion
This paper introduced the Set-Sequence model, an efficient architecture for capturing latent cross-sectional dependencies without manual feature engineering. By learning a shared cross-sectional summary at each period with a Set model and subsequently augmenting individual unit time series for a Sequence model, our approach demonstrates strong empirical performance on synthetic, stock return, and mortgage risk tasks, significantly outperforming benchmarks. While many existing multivariate models are powerful, they are often not natively designed to exploit unit exchangeability in multivariate financial time series. The Set-Sequence model offers a flexible and scalable solution.
Limitations: Future work should include a deeper exploration of alternative set architectures within the Set model, an investigation into the model’s performance on datasets with weaker exchangeability assumptions, and further analysis of the interpretability of the learned set summaries.
Appendix A Additional Literature Review

### v2

@@ 5.2 Generalization Across Number of Input Units
The task is to accurately predict state transitions when only a fraction $n$ of the total $M$ units are observed at inference time. This simulates having a large, evolving universe of loans but only observing a limited subset. We develop a near-optimal Kalman Filter baseline that assumes the true transition dynamics are known, with uncertainty only arising from observing a fraction of the total defaults. This oracle-like baseline represents a performance upper bound for the task. The full derivation is in Appendix E.3. During training we draw the number of units per batch as $D\sim(1-\gamma)\cdot\delta_{M}+\gamma\cdot\text{LogUnif}(1,M)$ with $\gamma=0.08$; thus most batches use all $M=1000$ units, while a few use smaller pools to encourage generalization. Further details are in Appendix E.
We evaluate the model’s ability to predict transitions to the absorbing (rare) state using AUC, correlation, and $R^{2}$. Formal definitions are in Appendix E.
@@
Figure 2: Comparison between the Set-Sequence model, MHA-Sequence model, and the Kalman Filter, by considering the AUC, $R^{2}$, and Correlation, each for the absorbing (rare) state. The Set-Sequence model reaches near the oracle-like Kalman Filter model performance across the full range of observed units for inference, for all the considered metrics.
Figure 2 shows that both the Set-Sequence and MHA-Seq models generalize well across all inference sizes, with performance approaching the oracle-like Kalman Filter baseline even when trained on a finite dataset. A single model effectively handles input sizes from 1 to 1000 units. While the MHA-Seq variant performs slightly better, it comes at a significant computational cost due to its quadratic scaling. The Set-Sequence model is interpretable. Table 3 shows a high correlation (up to 0.951) between a learned set summary and the true latent factor $\lambda_{0,t}$, with the correlation increasing as more units are observed. Appendix E.2 contains visualizations.
@@
Table 3: Average correlation for 100 samples, each with 100 time steps between the set summary at layer i and the true latent variable, $\lambda_{0,t}$. Here and in subsequent tables, best (second best) values in each column are in bold (underlined).
@@
6 Application: Equity Portfolio Construction
We evaluate the Set-Sequence model on a portfolio construction task. Given a universe of $N$ stocks, at the current time, the model outputs $N$ portfolio weights, indicating the portfolio to be held. The portfolio returns are realized from the inner product between the next day returns and the chosen portfolio weights. The goal is to optimize the Sharpe ratio, a measure of the risk-adjusted returns. This is a common objective in the asset pricing literature, as in Guijarro-Ordonez et al. (2022).
@@ 6 Application: Equity Portfolio Construction
We construct a dataset of 36,600 stocks (source: CRSP). For each period, we train on 8 years and test the model on the following year. For each model fitting, we filter out stocks where any returns in the train or test period are not available. We include only the 500 assets with the largest market cap at the last quarter of the training period, mimicking the S&P 500 index. This ensures that we consider assets that are liquid with small bid-ask spreads, see Guijarro-Ordonez et al. (2022). We use 79 features from Chen et al. (2024), applying cross-sectional rank normalization. They are based on recent cumulated returns, volatility, volume, and quarterly firm characteristics (such as book-to-market ratio). See Tables 10 and 11 in Appendix F.
@@
Let $X^{i}_{t}$ be the covariates at time $t$, and let $Y^{i}_{t}$ be the excess daily returns for asset $i$ (returns with the risk free rate subtracted), where $X_{t}$ is the covariate vector, and $Y_{t}$ is the return vector. We then use $X_{1},\dots,X_{t-1}$ to predict $\hat{w}_{t}$, the portfolio weights at the next step. Given a series of daily excess returns, $r_{1},\dots,r_{T}$, where $r_{t}=Y_{t}^{T}\hat{w}_{t}$, the daily Sharpe ratio is $SR=\bar{r}/\sqrt{\frac{1}{T}\sum_{t=1}^{T}(r_{t}-\bar{r})^{2}}$ where $\bar{r}=\frac{1}{T}\sum_{t=1}^{T}r_{t}$. We $L_{1}$ normalize the portfolio weights at each timestep to maintain unit leverage ($w_{t}||_{1}=1$). Following Guijarro-Ordonez et al. (2022) we optimize SR-SR.
Baselines, Metrics and Results
We compare with the sequence baselines in Section 4. Each baseline is trained on a single unit at a time with weights shared across units—this generalizes far better than feeding the entire cross-section, which performs poorly (see Section 5.1). We further compare with task-specific models: the CNN-Transformer and the baselines of Guijarro-Ordonez et al. (2022), which use aligned dataset, time span, and feature set (following Chen et al. (2024)).
The main metric to evaluate the performance is the annualized Sharpe ratio. We also report the percentage yearly return, the yearly return volatility, the average daily turnover, defined as the average $L_{1}$ distance between the weights in consecutive days. Beta is the covariance of portfolio and market returns divided by the market variance, measuring market exposure. A low beta is desirable to ensure returns are independent of market direction.
Table 4 shows the aggregate performance. The Set-Sequence model outperforms every sequence model baseline. We note a high annualized mean Sharpe ratio over the full period of 4.82, 22% higher than the second best sequence model (S4), and 32% higher than the LongConv model, which the Set-Sequence model uses for the sequence component in the experiment. Compared with the other sequence models, the Set-Sequence model shows robust results, with a Sharpe ratio standard deviation over 5 random initialized training runs of 0.12, compared with 0.52 for the Transformer and 0.29 for S4. Figure 8 in the Appendix shows the cumulative returns for the models.
Table 4: Summary statistics for the equities task out of sample (Jan. 2002–Dec. 2021). Each model is trained five times on different random seeds; all values are the mean over those runs, and the Sharpe ratio is reported as $\text{mean}\pm\text{std}$. The Sharpe Ratio, Mean Return, and Std-Dev Return are annualized. Beta is relative to the market, while Daily Turnover and Short Fraction are daily averages.
@@
LongConv $3.64\,\pm\,0.14$ 12.8 3.51 0.033 0.97 0.48
S4 $\underline{3.94}\,\pm\,0.29$ 13.5 3.43 0.028 0.90 0.48
H3 $3.46\,\pm\,0.28$ 10.6 3.08 0.026 1.12 0.48
Transformer $3.65\,\pm\,0.52$ 12.9 3.62 0.035 0.83 0.47
Hyena $2.91\,\pm\,0.33$ 8.7 3.02 0.032 1.28 0.47
Set-Sequence (Ours) $\textbf{4.82}\,\pm\,0.12$ 13.0 2.69 0.028 0.91 0.48
We also compare the Set-Sequence model with domain specific models on the time period Jan. 2002 - Dec. 2016 to be aligned with time period of prior work. Table 5 reports that the general Set-Sequence model outperforms the CNN-Transformer designed specifically for the equities task by 42 % in terms of the Sharpe ratio. In Section F in the Appendix we include an analysis of the Sharpe ratio when we account for transaction costs. We use a net Sharpe objective function and show that the Set-Sequence model still outperforms all baselines in the presence of transaction costs.
Table 5: Out-of-sample annualized performance metrics from 2002 to 2016 for the stock portfolio construction task. All CNN-Transformer models are based on the IPCA method described in Guijarro-Ordonez et al. (2022). For the Set-Sequence model the results are the mean over 5 training initialization seeds. For the Sharpe we also report its standard deviation over the seeds.
@@
Table 6: Comparing out-of-sample performance of different models on the mortgage risk prediction task on the test set. The average AUC is the mean of the AUCs to go between different states. The average transition probabilities are the average probability for the correct class over the full test set.
@@
Set-Seq (Ours) $\mathbf{0.200}$ $\mathbf{0.683}$
7 Application: Mortgage Risk Prediction
Next we evaluate the Set-Sequence model on an important mortgage risk prediction task, with the goal to predict the mortgage state (current, 30 days delinquent, 60 days delinquent, 90+ days delinquent, Foreclosure, Paid Off, and Real-Estate Owned) in the next month given a history of 52 covariates (FICO credit score, loan balance, current interest rate, etc., see Table 12 in the Appendix).
Dataset
We source the data from 4 ZIP codes in the greater LA area from CoreLogic; these are among the ZIP codes in the US with most active mortgages. We follow the data filtering procedure of Sadhwani et al. (2020), resulting in 5 million loan-month transitions from 117,523 loans. We use the following train, validation, test split: 1/1994 - 6/2009, 7/2009 - 12/2009, 1/2010 - 12/2022.
Baselines, Metrics and Results
We compare the Set-Sequence model with the domain specific models presented in Sadhwani et al. (2020), representing the current state of the art, along with the sequence models from Section 4. The models from Sadhwani et al. (2020) include a 5-layer neural network baseline with early stopping and dropout regularization, and a logistic regression baseline. Both baselines use the features at the current time to predict the state in the following month. For each training batch, for a given time we sample 2500 loans out of all active ones. We use a sequence length of 50, and use a multi-class cross entropy loss during training.
Table 6 shows that the Set-Sequence model achieves the best performance on cross-entropy loss and average AUC, outperforming all baselines. In particular, it improves average AUC by 4 points over the best domain-specific baseline (5-layer neural network). In Figure 3 we see that the Set-Sequence model robustly outperforms the prior state of the art model on the task, with better AUC for 22 out of 25 transitions, including the most common and economically important transitions, such as current to paid off. In Figure 12 in the Appendix, the foreclosure rate, a known source of cross unit dependency, is shown to be highly correlated with the set summary in the first Set-Sequence layer, indicating the interpretability of the learned set summaries.
Figure 3: AUC gain (Set-Sequence minus baseline Sadhwani et al. (2020)) for each transition. Deeper green indicates a larger Set-Sequence advantage; red indicates the baseline performed better. Larger circles indicate more common transitions, with size proportional to the log of their frequency. Only transitions occurring at least 10 times are included. In parentheses is the standard deviation over 10 random subsets of the evaluation set.
Appendix G supplements the discussion above with additional details and results. These include additional experiments for year-by-year performance and additional interpretability results.
@@ 8 Conclusion
This paper introduced the Set-Sequence model, an efficient architecture for capturing latent cross-sectional dependencies without manual feature engineering. By learning a shared cross-sectional summary at each period with a Set model and subsequently augmenting individual unit time series for a Sequence model, our approach demonstrates strong empirical performance on synthetic, investment, and risk prediction tasks, significantly outperforming benchmarks. While many existing multivariate models are powerful, they are often not natively designed to exploit unit exchangeability in multivariate time series. The Set-Sequence model offers a simple and scalable solution.
Appendix A Theory
A.1 Proof of Proposition 1 and a Corollary
Proof.
We count scalar operations to within constant factors and then sum over $t=1,\dots,T$.
Preliminaries.
At any time $t$, forming per-unit embeddings/fusions by dense linear maps on $x^{(i)}_{t}\in\mathbb{R}^{d}$ costs $\Theta(d)$ per unit, hence $\Theta(Md)$ across all units. Mean/sum pooling of $M$ vectors in $\mathbb{R}^{d}$ is a single pass over $Md$ scalars, i.e. $\Theta(Md)$.
For cross-sectional attention over $M$ tokens of width $d$, the standard forward pass per head entails: (i) Q/K/V linear maps: $\Theta(Md^{2})$, (ii) score matrix $QK^{\top}$: $\Theta(M^{2}d)$, (iii) value aggregation $({\rm softmax}(QK^{\top}))V$: $\Theta(M^{2}d)$, (iv) output projection: $\Theta(Md^{2})$, and an FFN of width $\Theta(d)$ per token: $\Theta(Md^{2})$. Grouping linear terms yields $\Theta(Md^{2})$ and the pairwise terms yield $\Theta(M^{2}d)$. Multi-head with a constant number of heads only changes constants.
When each scalar becomes a token (Naïve feature-as-token), there are $N=Md$ tokens. With width $\Theta(1)$ per token, Q/K/V/out projections are $\Theta(N)$ and attention multiplies are $\Theta(N^{2})=\Theta((Md)^{2})$, which dominates.
(i) Set–Seq.
At each $t$, compute per-unit transforms and fuse with the pooled summary. This is one linear scan over all $M$ units with $d$ features: $\Theta(Md)$ to embed, $\Theta(Md)$ to pool, and $\Theta(Md)$ to fuse; thus $\Theta(Md)$ up to constants. Summed over $T$ time steps gives $\Theta(TMd)$ for the cross-sectional part. Then we run one temporal model per unit on width $d$, for a total of $\Theta\!\big(M\,C_{\mathrm{seq}}(T,d)\big)$. Hence (i).
(ii) Attention–Seq (M tokens, width d).
Per $t$, cross-sectional attention over $M$ tokens of width $d$ costs $\Theta(M^{2}d)$ for scores/aggregation and $\Theta(Md^{2})$ for projections/FFN, as detailed above. Summed over $T$ time steps gives $\Theta\!\big(T[M^{2}d+Md^{2}]\big)$. Temporal processing is still per unit at width $d$, adding $\Theta\!\big(M\,C_{\mathrm{seq}}(T,d)\big)$. Hence (ii).
(iii) Naïve feature-as-token (N=Md).
Per $t$, there are $N=Md$ tokens. With token width $\Theta(1)$, the attention-matrix multiplies (scores and aggregation) cost $\Theta(N^{2})=\Theta((Md)^{2})$ and dominate the $\Theta(N)$ projection costs. Summed over $T$ gives $\Theta(T(Md)^{2})$. Temporal processing remains per unit at width $d$, adding $\Theta\!\big(M\,C_{\mathrm{seq}}(T,d)\big)$. Hence (iii).
(iv) Gated-Selection–Seq.
By assumption, the cross-sectional gating/selection matrix is computed once (not per time) from $M$ unit summaries of width $d$ via a dense similarity, which is $\Theta(M^{2}d)$: there are $\Theta(M^{2})$ pairs and computing each affinity costs $\Theta(d)$. Thereafter, at each time step the layer behaves as Set–Seq (no pairwise recomputation), costing $\Theta(Md)$ per $t$, hence $\Theta(TMd)$ across the sequence, plus the per-unit temporal total $\Theta\!\big(M\,C_{\mathrm{seq}}(T,d)\big)$. Hence (iv).
(v) Full-Stacked Temporal.
At each $t$, stack all $M$ unit features into a single vector of width $Md$ (reshaping costs $O(Md)$ and is dominated by the temporal block). Apply a single temporal model of width $Md$ over length $T$, whose cost is by definition $C_{\mathrm{seq}}(T,Md)$. No per-unit temporal pass occurs. Hence (v).
Conclusion.
Combining the per-time costs (and the one-off gate in (iv)) with the temporal costs yields the stated asymptotics. Backpropagation traverses the same computations with constant-factor overhead, leaving the $\Theta(\cdot)$ orders unchanged. This completes the proof. ∎
Corollary 1 (Plugging in standard temporal self-attention).
If the temporal model is standard self-attention with cost $C_{\mathrm{attn}}(T,w)=\Theta(T^{2}w)+\Theta(Tw^{2})$, then:
(i) Set–Seq: $\quad\Theta(TMd)\;+\;\Theta\!\big(M[T^{2}d+Td^{2}]\big),$
(ii) X-section Attn (M tokens): $\quad\Theta\!\big(T[M^{2}d+Md^{2}]\big)\;+\;\Theta\!\big(M[T^{2}d+Td^{2}]\big),$
(iii) Naïve (Md tokens): $\quad\Theta\!\big(T(Md)^{2}\big)\;+\;\Theta\!\big(M[T^{2}d+Td^{2}]\big),$
(iv) Gated-Selection–Seq: $\quad\Theta(M^{2}d)\;+\;\Theta(TMd)\;+\;\Theta\!\big(M[T^{2}d+Td^{2}]\big),$
(v) Full-Stacked Temporal: $\quad\Theta\!\big(T^{2}Md+T(Md)^{2}\big).$
A.2 Proof of Proposition 2
Proof.
Step 1 (Empirical moments as pooled monomials). For a multi-index $\alpha=(\alpha_{1},\dots,\alpha_{d})\in\mathbb{N}_{0}^{d}$ with $|\alpha|=\sum_{j=1}^{d}\alpha_{j}$, define $x^{\alpha}=\prod_{j=1}^{d}x_{j}^{\alpha_{j}}$ and, for any $(x^{(1)},\dots,x^{(M)})\in K^{M}$, the empirical moment
$m_{\alpha}=\frac{1}{M}\sum_{i=1}^{M}(x^{(i)})^{\alpha}\;=\;\frac{1}{M}\sum_{i=1}^{M}\prod_{j=1}^{d}(x^{(i)}_{j})^{\alpha_{j}}.$
Let $\phi_{k}(x)$ stack all monomials with $|\alpha|\leq k$. Then $\frac{1}{M}\sum_{i=1}^{M}\phi_{k}(x^{(i)})$ stacks all $\{m_{\alpha}:|\alpha|\leq k\}$. The number of such monomials is $\sum_{r=0}^{k}\binom{d+r-1}{r}=\binom{d+k}{k}$ (stars-and-bars).
Step 2 (Factorization through the empirical measure; well-defined and continuous). Define the empirical measure $\mu=\frac{1}{M}\sum_{i=1}^{M}\delta_{x^{(i)}}$. Any permutation of $(x^{(1)},\dots,x^{(M)})$ leaves $\mu$ unchanged and, by invariance, leaves $G$ unchanged. Define
$\widetilde{G}(\mu)\;:=\;G(x^{(1)},\dots,x^{(M)})\quad\text{for any ordering that realizes }\mu,$
which is well-defined because equal empirical measures correspond to permutations of the same multiset and $G$ takes equal values on such permutations. The map $(x^{(1)},\dots,x^{(M)})\mapsto\mu$ is continuous in the weak topology on $K$ (if $x^{(i)}\to y^{(i)}$, then for every continuous $f$ on $K$, $\int f\,d\mu\to\int f\,d\nu$ where $\nu=\tfrac{1}{M}\sum_{i}\delta_{y^{(i)}}$). Since $G=\widetilde{G}\circ\mu$ and $G$ is continuous on $K^{M}$, $\widetilde{G}$ is continuous on the compact image of $K^{M}$ under this map.
Step 3 (Polynomial moments separate empirical measures). On compact $K$, polynomials are dense in $C(K)$ (Stone–Weierstrass). Thus for any two distinct empirical measures $\mu\neq\nu$ there exists a polynomial $p$ with $\int p\,d\mu\neq\int p\,d\nu$. Hence the collection $\{\int p(x)\,d\mu(x):p\text{ polynomial}\}$ separates empirical measures.
Step 4 (Approximate $\widetilde{G}$ by a polynomial in finitely many moments). The domain of $\widetilde{G}$ is compact, and the algebra generated by polynomial moments contains constants and separates points. By Stone–Weierstrass, for any $\varepsilon>0$ there exists a polynomial $\Psi$ in finitely many moments $\{\int p_{\ell}\,d\mu\}_{\ell=1}^{L}$ such that
$\sup_{(x^{(1)},\dots,x^{(M)})\in K^{M}}\big\|\Psi(\mu)-\widetilde{G}(\mu)\big\|_{\infty}\;<\;\varepsilon/2.$
Each $\int p_{\ell}\,d\mu$ equals $\frac{1}{M}\sum_{i=1}^{M}p_{\ell}(x^{(i)})$, and for some degree $k$ each $p_{\ell}$ is a linear combination of monomials of total degree $\leq k$.
Step 5 (Realize $\Psi$ via pooled monomials). With $\phi_{k}$ as in Step 1, there exists a (multivariate polynomial) map $\rho:\mathbb{R}^{\binom{d+k}{k}}\to\mathbb{R}^{q}$ such that
$\Psi(\mu)\;=\;\rho\!\Big(\frac{1}{M}\sum_{i=1}^{M}\phi_{k}(x^{(i)})\Big)\quad\text{for all }(x^{(1)},\dots,x^{(M)})\in K^{M}.$
Step 6 (Uniform error bound). Combining Steps 4–5,
$\sup_{(x^{(1)},\dots,x^{(M)})\in K^{M}}\Big\|\rho\!\Big(\tfrac{1}{M}\sum_{i=1}^{M}\phi_{k}(x^{(i)})\Big)-G(x^{(1)},\dots,x^{(M)})\Big\|_{\infty}\;\leq\;\varepsilon/2\;<\;\varepsilon,$
after tightening constants if desired. This completes the proof. ∎
Remark (Quadratic case).
Taking $k=2$ and $\phi_{2}(x)=[\,x,\ \mathrm{vec}_{\mathrm{sym}}(xx^{\top})\,]$ recovers first and second empirical moments exactly, so any permutation-invariant quadratic statistic of the cross-section is obtained by a linear $\rho$.
Appendix B Additional Literature Review